# Can Charge Transfer Across C–H···O Hydrogen Bonds Stabilize Oil Droplets in Water?

### v1

@@ Abstract
Oil-water emulsions resist aggregation due to the presence of negative charges at their surface that leads to mutual repulsion between droplets, but the molecular origin of oil charge is currently under debate. While much evidence has suggested that hydroxide ions must accumulate at the interface, an alternative perspective attributes the negative charge on the oil droplet to not an ionic species, but charge transfer of electron density from water to oil molecules. While the charge transfer mechanism is consistent with the correct sign of oil charge based on electrophoresis experiments, it is just as important to provide good estimates of the magnitude of the negative charge to explain emulsion stability. Here we show using energy decomposition analysis that the amount of net flow of charge from water to oil is negligibly small due to nearly equal forward and backward charge transfer through weak oil-water interactions, such that oil droplets would be unstable and coalesce, contrary to experiment. The lack of charge transfer also explains why vibrational sum frequency scattering reports a blue shift in the oil C-H frequency when forming emulsions with water, which arises from Pauli repulsion due to localized confinement at the interface.
@@ 1 Introduction
Oil and water are capable of forming stable emulsions in which the minor component forms droplets dispersed throughout the solvent. Oil-water emulsions are thought to be electrostatically stabilized by surface charges that lead to mutual repulsion between droplets to confer stability. For the case of oil droplets in water, electrophoresis experiments consistently show that the oil droplets migrate toward the anode under an applied electric field1, 2, which has typically been interpreted to mean that the oil droplets preferentially adsorb negative ions present in the aqueous solution. This gives rise to a stationary layer of ions that adhere to the oil particle and travel with the droplet under the applied field, creating an electrostatic potential difference with the mobile fluid at the “slip plane” which defines the $\zeta$ potential. The $\zeta$ potential can be inferred from the electrophoretic mobility of the droplet and is frequently used as a gauge of emulsion stability3, 4, with $\zeta$ potential magnitudes less than $\sim$30 mV resulting in flocculation or coalescence5. For oil-in-water emulsions prepared at neutral pH, the $\zeta$ potential is typically around --60 to --90 mV,6, 7, 8 indicating excellent long term droplet stability.
@@
Because stable oil-in-water emulsions can be prepared without added ions, it is frequently assumed that hydroxide ions from the auto-dissociation of water preferentially adsorb to the water oil-interface.9, 10 Indeed, the $\zeta$ potential varies strongly with the pH, including going through an isoelectric point around pH = 3.1 In 2004, Beattie and Djerdjev estimated the surface charge density of interfacial ions by combining electroacoustic droplet sizing with measurements of the NaOH required to maintain a constant pH during homogenization.11 They observed that the amount of OH- added increased linearly with increasing interfacial area, suggesting an intrinsic and constant surface charge density at the oil-water interface. Under the assumption of excess hydroxide ions they estimated the surface charge to be --0.28 e-/nm2. More recently using second harmonic generation (SHG) experiments, Gan et al. found that hydroxide, not hydronium, showed preferential absorption at the surface of an hexadecane emulsion in water.10 Their estimates of hydroxide surface charge of --0.14 e-/nm2, which is comparable to the value obtained previously by Beattie and Djerdjev, would support an adsorption free energy of $\sim$-∼ -8.3 kcal/mol.10
@@
However a different perspective has been offered over the last 15 years that attributes the negative charge on the oil droplet not to ionic species, but an accumulation of excess electron density at the interface arising from charge transfer (CT).6 In 2011 Vacha and co-workers6 combined mobility measurements, sum frequency scattering (SFS) experiments, and molecular dynamics simulations. The mobility measurements measured a $\zeta$ potential of --81 mV, while the SFS experiments showed no selective adsorption of hydroxide ions to the oil-water interface. The molecular dynamics (MD) simulations found that a small amount of charge was transferred between water hydrogen bonds, not between water and oil, using the TIP4P-FQ water model that is capable of modeling CT.12 They calculated a water surface charge density of $-0.0001$ to $-0.0028$ e-/nm2, and subsequent studies employing similar charge equilibration models13, 14 also reported comparable surface charge densities in the range of $-0.001$ to $-0.002$ e-/nm2. However, the water surface charge estimates discussed in these studies primarily refer to CT interactions between water molecules rather than CT between water and oil.6, 13, 14.
@@
In subsequent work Pullanchery and co-workers introduced vibrational sum frequency scattering (VSFS) measurements on oil-water emulsions15, and found spectroscopic signatures in which the oil C–H modes blue shift while the water vibrations red shift at the interface. This spectroscopic signature was claimed to be supportive for a CT mechanism, since they again found no experimental evidence from VSFS of interfacial OH radicals or OH- species.15 Supporting theoretical calculations further proposed that CT occurs across C–H···O hydrogen bonds15. The CT mechanism has also been analyzed from theoretical calculations using the density-derived electrostatic and chemical (DDEC)16 atomic charge partitioning scheme, which found a charge of --0.015 e-/nm2, but only after accumulation into the droplet and deep in the oil phase.17 This is an order of magnitude higher than previous CT calculations.6, 13, 14 More recently in 2024, Roke and co-workers found that while oil mobility changes as a function of increasing pH, the blue-shifted vibrational signatures of the C–H peak remain unperturbed by increasing pH which they claim confirms a CT mechanism18. To explain the mobility results, they conducted simulations and found that under a 0.1 V/ Å electric field oriented along the z-axis, neopentane exhibited a net drift relative to the surrounding water which they attributed not solely to CT but also to field-induced ’dynamic polarization’, to help explain the sensitivity of mobility, but not vibrations, to pH.18
@@
While the CT offers an intriguing explanation for interfacial charging of the oil phase, there are several reasons to doubt its viability for explaining stability, electrokinetic experiments, and spectroscopic measurements of oil-water emulsions. To address these concerns, we employ model cluster calculations as well as a polarizable force field simulation of an extended water–hexane interface to quantify the amount of CT using quantum mechanical energy decomposition analysis based on the absolutely localized molecular orbital (ALMO-EDA) method19, 20, 21, 22, 23. Our findings indicate that the surface charge densities computed by ALMO-EDA are one to two orders of magnitude smaller than the DDEC atomic charge partitioning scheme or charges obtained from Mulliken populations. Furthermore, true CT analysis shows any CT from water to oil is largely negated by back-transfer of electron density24 from oil to water, in which back-transfer is a natural energy lowering process when one correctly accounts for electron flow. Therefore such small amounts of net CT can’t possibly explain oil-water emulsion stability and hence does not explain oil charging and electrophoresis experiments. Finally, we reaffirm again that vibrational blue shifts of C-H actually arise from Pauli repulsion whereas CT would redshift these same modes25, 26, 22, and thus a CT mechanism can’t be the origin of the VSFS spectroscopic observations. Finally we show that explanations based on dynamic polarization to explain mobility18 lack evidence due to simulation errors that fail to exhibit directional drift under an applied electric field.
@@ 2 Results
Beginning with accurate state-of-the-art density functional theory (DFT) calculations27, the ALMO-EDA framework19, 20, 21, 22, 23 decomposes the overall intermolecular interaction energy, $\Delta E_{\text{INT}}$ between several species or fragments into three physically interpretable contributions, providing insight into their relative significance.
@@
Within the ALMO-EDA framework, electron density rearrangement associated with CT can be interpreted as donor-acceptor orbital interactions. CT occurs when an occupied orbital on one molecule donates electron density to a complementary virtual orbital on another molecule. This interaction is quantified through complementary occupied-virtual pair (COVP)19, 28, 29 analysis. In what follows we consider a single dominant forward COVP and a single dominant backward COVP, because they are sufficient to capture approximately 75% of the overall CT contribution to both the energy and net charge, making them a compact summary of the CT orbital picture. A more detailed description of the ALMO-EDA method, as implemented in Q-Chem30, and further details of forward and backward CT using COVP analysis are provided in the Methods section.
@@ 2.1 Charge transfer across the C-H···O hydrogen bond
We begin with the prototype system for understanding CT based on the strong hydrogen bond O-H···O formed by the the water dimer, in order to compare it to C-H···O interactions in which it is claimed that CT occurs from water to oil15. As seen in Figure 1A, the COVP indicates that hydrogen bonding in the water dimer is composed of CT from an electron pair localized on an oxygen atom donating electrons to the O–H $\sigma$ anti-bonding orbital on the other water. For the water dimer, CT occurs predominantly in the forward direction, from the hydrogen-bond acceptor water molecule to the donor water molecule, with a net CT of 2.67 me- quantified in full by ALMO-EDA.
To accurately assess CT in oil–water interactions, we next analyze the C-H···O interaction that has been widely recognized as another, albeit weaker, type of hydrogen bond.25, 31 Many studies impose a linear constraint when analyzing C-H···O interactions25, however their relatively weak strength makes them more susceptible to bending and less sensitive to deviations from their equilibrium bond length25. This is seen in the optimized structures of a hexane molecular fragment (C6H14) interacting with one water molecule as shown in Figure 1B, in which the optimized structure exhibits a C-H···O bond angle of approximately 120°, with the hydrogens in water positioned closer to hexane than the oxygen atoms. This significantly deviates from the linear hydrogen bond configuration observed in the water dimer or idealized C-H···O configurations.
@@
Figure 1: Charge transfer analysis for optimized water–hexane complexes compared to the water dimer. Optimized structures and dominant COVPs in forward and backward charge transfer processes of A (H2O)(H2O), B (C6H14)(H2O), and C(C6H14)(H2O1—H2O2) complexes. Forward CT refers to H2O $\rightarrow$ C6H14, and backward CT refers to C6H14 $\rightarrow$ H2O. D and E represent charge transfer and ALMO-EDA energy analysis of optimized complexes in A, B, C, and C1 and C2 that represent the complexes (C6H14)(H2O1) and (C6H14)(H2O2) after removing one water molecule from complex C. All structural optimizations were performed at the $\omega$B97X-V/def2-TZVP level, and AMLO-EDA calculations were conducted using $\omega$B97X-V/def2-TZVPD. CT unit me-. The original data for D and E are provided in Supplementary Tables S1 and S2.
In Figure 1C we see that the presence of a second water molecule enhances the forward CT by a factor of $\sim$1.8 compared to the hexane-water dimer in Figure 1B. This enhancement arises from the formation of a stronger electron donor through hydrogen bonding with the second water molecule. However, the increased forward CT is accompanied by an increase in backward CT, as the second water molecule becomes more electron-deficient, with the virtual orbitals in the backward CT primarily localized on the second water molecule. In this case the C-H group returns electron density to the anti-bonding orbital of the second water given the close range of H··H interactions. To further investigate the role of the C-H···O hydrogen bond geometry in the CT mechanism, we remove one water molecule from the (C6H14) (H2O1—H2O2) to create (C6H14)(H2O1) (C1) and removing the other water molecule to create (C6H14)(H2O2) (C2). As seen in Figure 1C, C1 has a well-defined hydrogen bond with a C··O distance of $\sim$3.35Å, resulting in a net CT of 0.36 me-, whereas C2 lacks an optimal hydrogen bond configuration, leading to a net CT in the reverse direction from hexane to water of --0.28 me-, which explains the negligible net charge flow in Figure 1C. These findings indicate that even though the presence of an additional water molecule enhances forward CT, the overall net electron transfer remains negligibly small, and furthermore it accumulates on water not hexane.
@@ 2.2 Charge transfer at an extended oil-water interface
Analysis of minimized dimer and trimer configurations reveal that forward and backward CT between water and hexane nearly cancels. However, the CT mechanism at the extended water–oil interface may be significantly more intricate due to thermal fluctuations and a more extended network of hydrogen-bond interactions between waters and hydrophobic association of hexanes. To gain further insight into the CT process at an extended water–oil interface, molecular dynamics at ambient conditions using the polarizable AMOEBA force field32, 33 was employed; simulation details are provided in the Methods section.
@@
From the MD simulations we draw two types of interfacial hexane-water clusters. Figure 2A is a representative snapshot of the 200 clusters in which the water molecules are localized around multiple hexanes, whereas Figure 2B is a representative snapshot of the 50 clusters in which interfacial water molecules form more extensive hydrogen-bonded network around the hexane. From these clusters we derive the probability density distribution in Figure 2C that shows a high-probability region near a CH…O angle of 120° for C-O distances $>$ 3.5 Å, consistent with the C-H···O configurations of the ab initio optimized clusters in Figure 1B,C. The broad distribution means that the C-H···O hydrogen bond is significantly weaker and more distorted compared to the highly directional hydrogen bonds found for even thermalized water–water interactions which experience significant CT. This would indicate that CT in the forward direction from water to oil will be much smaller at an extended oil-water interface as well. In Figure 2D, the distribution of distances between hexane and water hydrogens is centered at 1.85 Å, whereas the distribution of water oxygen to hexane hydrogen distances is peaked at 2.2 Å, and the two distributions exhibit minimal overlap. This interfacial organization implies that the hydrogen of surface water orientated towards the oil (“dangling O-H groups”) will be a conduit for back-transfer of electron density from oil to water.
@@
This is borne out in Figure 2E which illustrates the ensemble nature of CT dynamics at the oil-water interface. The forward CT distribution of the interfacial clusters peaks at 2.09 me-/nm2, which arises from the C-H···O hydrogen bond at the interface, and similar to the small clusters in Figure 1. But again this is balanced by electron backflow from the CT distribution centered around -1.44 me-/nm2 due to the presence of short-range H···H
@@
Figure 2: Water/oil interface structures. (A, B) Sample representation of the clusters considered for the ALMO-EDA calculations highlighting the spatial arrangement of water molecules (red and white spheres) around hexane molecules (cyan). The extended hydrogen-bonded water network at the interface is illustrated in pink. (C) Probability density distribution illustrating the spatial and angular relationship between water oxygen and hexane. The x-axis represents the C–O distance, while the y-axis shows the C-H···O angle, describing the relative orientation of water molecules with respect to hexane’s C–H bonds. (D) Probability distribution of the closest distance between hydrogen atoms in hexane with water hydrogens and oxygens. This is consistent with the radial distribution functions (Supplementary Figure S1), where the H···H intermolecular distribution is more short-ranged than the O···H distribution in the water–hexane system. (E) Probability density distributions of Forward CT, Backward CT, and Net CT. Solid lines for direct water–hexane interactions clusters in (A) and dashed lines correspond to the clusters in (B). The original data is provided in Supplementary Table S3 and S4. (F) Probability density distributions of water oil inter-molecule interaction energy components. The legend includes the percentage contribution of each energy component relative to the total interaction energy. Forward CT refers to H2O $\to$ C6H14, and Backward CT refers to C6H14 $\to$ H2O. Cluster selection details are provided in the Methods section.
interactions at the interface that defines the source of backward CT as also seen in Figure 1. The average net CT of 0.00065 e-/nm2 (0.65 me-/nm2) obtained from ALMO-EDA corresponds to 0.00007 e-/hydrogen bond, closely matching the 0.0001 e-/hydrogen bond observed in the hexane–water clusters in Figure 1.
@@
Furthermore these conclusions are consistent between the clusters with localized waters near hexane or the clusters incorporating the extended hydrogen bonding network since interfacial CT is predominantly dictated by direct water–hexane interactions, whereas the broader hydrogen bonding network within the interfacial region exerts only a subtle perturbative influence (average net CT of 0.00061 e-/nm2). Finally Figure 2F shows that the frozen energy term constitutes the dominant component (61.6%) to water-hexane interactions, with similar percentages of polarization energy and CT making up the remaining energetic interactions. The estimated 20.3% of CT energy between water and hexane is a proportion similar to that observed in Figure 1 for small water–hexane clusters B and C, and is energetically the manifestation of electron delocalization in stabilizing the system.
2.3 Surface charge and oil-water emulsion stability

### v3

@@ Abstract
Oil-water emulsions resist aggregation due to the presence of negative charges at their surface that leads to mutual repulsion between droplets, but the molecular origin of oil charge is currently under debate. While much evidence has suggested that ionic species must accumulate at the interface, an alternative perspective attributes the negative charge on the oil droplet to charge transfer of electron density from water to oil molecules. While the charge transfer mechanism is consistent with the correct sign of oil charge, it is just as important to provide good estimates of the charge magnitude to explain emulsion stability and electrophoresis experiments. Here we show using energy decomposition analysis that the amount of net flow of charge from water to oil is negligibly small due to nearly equal forward and backward charge transfer through weak oil-water interactions, such that oil droplets would be unstable and coalesce, contrary to experiment. The lack of charge transfer also explains why vibrational sum frequency scattering reports a blue shift in the oil C-H frequency when forming emulsions with water, which arises from Pauli repulsion due to localized confinement at the interface. Finally, unlike ions, neither charge transfer nor dynamic polarization can produce a finite conductivity needed to couple to electric fields that would explain electrophoretic mobility.
@@ 1 Introduction
Oil and water are capable of forming long-lived metastable emulsions in which the minor component forms droplets dispersed throughout the solvent. Oil-water emulsions are thought to be electrostatically stabilized by surface charges that lead to mutual repulsion between droplets to confer metastability. For the case of oil droplets in water, electrophoresis experiments consistently show that the oil droplets migrate toward the anode under an applied electric fieldcarruthers1938, hunter2013zeta, indicating that the oil droplets preferentially adsorb negative ionic species present in the aqueous solution. This gives rise to a stationary layer of ions that adhere to the oil particle and travel with the droplet under the applied field, creating an electrostatic potential difference with the mobile fluid at the “slip plane” which defines the $\zeta$ potential. The $\zeta$ potential can be inferred from the electrophoretic mobility of the droplet and is frequently used as a gauge of emulsion stabilityanderson1989colloid, ohshima1994simple, with $\zeta$ potential magnitudes less than $\sim$30 mV resulting in flocculation or coalescencekumar2017methods. For oil-in-water emulsions prepared at neutral pH, the $\zeta$ potential is typically around -60 to -90 mV,Vacha2011, agmon2016protons, Shi2025 indicating excellent longer term droplet metastability.
@@
Because stable oil-in-water emulsions can be prepared without added ions, it is frequently assumed that hydroxide ions from the auto-dissociation of water preferentially adsorb to the water oil-interface.beattie2009surface, Gan2017 Indeed, the $\zeta$ potential varies strongly with the pH, including going through an isoelectric point around pH = 3.carruthers1938 In 2004, Beattie and Djerdjev estimated the surface charge density of interfacial ions by combining electroacoustic droplet sizing with measurements of the NaOH required to maintain a constant pH during homogenization.Beattie2004 They observed that the amount of OH- added increased linearly with increasing interfacial area, suggesting an intrinsic and constant surface charge density at the oil-water interface. Under the assumption of excess OH- ions they estimated the surface charge to be -0.28 e-/nm2. More recently using second harmonic generation (SHG) experiments, Gan et al. found that hydroxide, not hydronium, showed preferential absorption at the surface of an hexadecane emulsion in water.Gan2017 Their estimates of OH- surface charge of -0.14 e-/nm2, which is comparable to the value obtained previously by Beattie and Djerdjev, would support an adsorption free energy of $\sim$-8.3 kcal/mol.Gan2017 Alternatively, some studies have suggested that the interfacial ions originate from trace amounts of surface-active impurities, such as amphiphilic molecules with carboxylic acid groups, which adsorb at hydrophobic interfaces and account for the negative zeta potential and its pH dependence.Roger2012a Netz and co-workers further demonstrated that including nanomolar concentrations of weakly acidic and basic impurities could reproduce experimental zeta potentials in theoretical models.uematsu2019impurity, uematsu2020nanomolar
@@
However a different perspective has been offered over the last 15 years that attributes the negative charge on the oil droplet not to ionic species, but an accumulation of excess electron density at the interface arising from charge transfer (CT).Vacha2011 In 2011 Vacha and co-workersVacha2011 combined mobility measurements, sum frequency scattering (SFS) experiments, and molecular dynamics simulations. The mobility measurements measured a $\zeta$ potential of -81 mV, while the SFS experiments showed no selective adsorption of OH- ions to the oil-water interface. The molecular dynamics (MD) simulations found that a small amount of charge was transferred between water hydrogen bonds, not between water and oil, using the TIP4P-FQ water model that is capable of modeling CT.lee2011effects They calculated a water surface charge density of 0.0001-0.0001 to 0.0028-0.0028 e-/nm2, and subsequent studies employing similar charge equilibration modelswick2012intermolecular, vacha2012charge also reported comparable surface charge densities in the range of 0.001-0.001 to 0.002-0.002 e-/nm2. However, the surface charge estimates from these studies primarily refer to CT interactions between water molecules rather than CT between water and oil, and furthermore are too small in magnitude to support metastability.Vacha2011, wick2012intermolecular, vacha2012charge.
@@
In subsequent work Pullanchery and co-workers introduced vibrational sum frequency scattering (VSFS) measurements on oil-water emulsionspullanchery2021charge, and found spectroscopic signatures in which the oil C–H modes blue shift while the water vibrations red shift at the interface. This spectroscopic signature was claimed to be supportive for a CT mechanism, since they again found no experimental evidence from VSFS of interfacial OH radicals, OH- ions, or contaminant species.pullanchery2021charge Supporting theoretical calculations further proposed that CT occurs across C–H···O hydrogen bondspullanchery2021charge. The CT mechanism has also been analyzed from theoretical calculations using the density-derived electrostatic and chemical (DDEC)manz2012improved atomic charge partitioning scheme, which found a charge of -0.015 e-/nm2, but only after accumulation into the droplet and deep in the oil phase.poli2020charge This is an order of magnitude higher than previous charge equilibration calculations that estimated CT surface density .Vacha2011, wick2012intermolecular, vacha2012charge More recently in 2024, Roke and co-workers found that while oil mobility changes as a function of increasing pH, the blue-shifted vibrational signatures of the C–H peak remain unperturbed by increasing pH which they claim confirms a CT mechanismPullanchery2024. To explain the mobility results, they conducted ab initio molecular dynamics (AIMD) simulations and found that under a 0.1 V/ Å electric field oriented along the z-axis, a single neopentane molecule exhibited a net drift relative to the surrounding water which they attributed not solely to CT but also to field-induced ’dynamic polarization’, to help explain the sensitivity of mobility, but not vibrations, to pH.Pullanchery2024
@@
While the CT offers an intriguing explanation for interfacial charging of the oil phase, there are several reasons to doubt its viability for explaining stability, electrokinetic experiments, and spectroscopic measurements of oil-water emulsions. To address these concerns, we employ model cluster calculations as well as a polarizable force field simulation of an extended water–hexane interface to quantify the amount of CT using quantum mechanical energy decomposition analysis based on the absolutely localized molecular orbital (ALMO-EDA) methodkhaliullin2008analysis, horn2016probing, Mao2018, Mao2019, Mao2021. Our findings indicate that the surface charge densities computed by ALMO-EDA are one to two orders of magnitude smaller than the DDEC atomic charge partitioning scheme or charges obtained from Mulliken populations. Furthermore, true CT analysis shows any CT from water to oil is largely negated by back-transfer of electron densityloipersberger2020variational from oil to water, in which back-transfer is a natural energy lowering quantum mechanical process when one correctly accounts for electron flow. Therefore such small amounts of net CT can’t possibly explain oil-water emulsion stability and hence does not explain oil charging and electrophoresis experiments. We reaffirm again that vibrational blue shifts of C-H actually arise from Pauli repulsion whereas CT would redshift these same modesgu1999fundamental, li2002physical, Mao2019, and thus a CT mechanism can’t be the origin of the VSFS spectroscopic observations. Finally we show that explanations based on dynamic polarization to explain mobilityPullanchery2024 lack evidence due to AIMD simulation errors that fail to exhibit directional drift under an applied electric field. This is important for the reason that neither the small density from CT nor the polarization fluctuations can sustain the conductivity needed to couple into the laterally applied electric fields to explain electrophoresis experiments.
@@ 2 Results
Beginning with accurate state-of-the-art density functional theory (DFT) calculationsMardirossian2017, the ALMO-EDA frameworkkhaliullin2008analysis, horn2016probing, Mao2018, Mao2019, Mao2021 decomposes the overall intermolecular interaction energy, $\Delta E_{\text{INT}}$ between several species or fragments into three physically interpretable contributions, providing insight into their relative significance.
@@
Within the ALMO-EDA framework, electron density rearrangement associated with CT can be interpreted as donor-acceptor orbital interactions. CT occurs when an occupied orbital on one molecule donates electron density to a complementary virtual orbital on another molecule. This interaction is quantified through complementary occupied-virtual pair (COVP)khaliullin2008analysis, veccham2021non, shen2022generalization analysis. In what follows we consider a single dominant forward COVP and a single dominant backward COVP, because they are sufficient to capture approximately 75% of the overall CT contribution to both the energy and net charge, making them a compact summary of the CT orbital picture. A more detailed description of the ALMO-EDA method, as implemented in Q-ChemEpifanovsky2021, and further details of forward and backward CT using COVP analysis are provided in the Methods section.
@@ 2.1 Charge transfer across the C-H···O hydrogen bond
We begin with the prototype system for understanding CT based on the strong hydrogen bond O-H···O formed by the the water dimer, in order to compare it to C-H···O interactions in which it is claimed that CT occurs from water to oilpullanchery2021charge. As seen in Figure 1A, the COVP indicates that hydrogen bonding in the water dimer is composed of CT from an electron pair localized on an oxygen atom donating electrons to the O–H $\sigma$ anti-bonding orbital on the other water. For the water dimer, CT occurs predominantly in the forward direction, from the hydrogen-bond acceptor water molecule to the donor water molecule, with a net CT of 2.67 me- quantified in full by ALMO-EDA.
To accurately assess CT in oil–water interactions, we next analyze the C-H···O interaction that has been widely recognized as another, albeit weaker, type of hydrogen bond.gu1999fundamental, horowitz2012carbon Many studies impose a linear constraint when analyzing C-H···O interactionsgu1999fundamental, however their relatively weak strength makes them more susceptible to bending and less sensitive to deviations from their equilibrium bond lengthgu1999fundamental. This is seen in the optimized structures of a hexane molecular fragment (C6H14) interacting with one water molecule as shown in Figure 1B, in which the optimized structure exhibits a C-H···O bond angle of approximately 120°, with the hydrogens in water positioned closer to hexane than the oxygen atoms. This significantly deviates from the linear hydrogen bond configuration observed in the water dimer or idealized C-H···O configurations.
@@
[…]
In Figure 1C we see that the presence of a second water molecule enhances the forward CT by a factor of $\sim$1.8 compared to the hexane-water dimer in Figure 1B. This enhancement arises from the formation of a stronger electron donor through hydrogen bonding with the second water molecule. However, the increased forward CT is accompanied by an increase in backward CT, as the second water molecule becomes more electron-deficient, with the virtual orbitals in the backward CT primarily localized on the second water molecule. In this case the C-H group returns electron density to the anti-bonding orbital of the second water given the close range of H··H interactions. To further investigate the role of the C-H···O hydrogen bond geometry in the CT mechanism, we remove one water molecule from the (C6H14) (H2O1—H2O2) to create (C6H14)(H2O1) (C1) and removing the other water molecule to create (C6H14)(H2O2) (C2). As seen in Figure 1C, C1 has a well-defined hydrogen bond with a C··O distance of $\sim$3.35Å, resulting in a net CT of 0.36 me-, whereas C2 lacks an optimal hydrogen bond configuration, leading to a net CT in the reverse direction from hexane to water of -0.28 me-, which explains the negligible net charge flow in Figure 1C. These findings indicate that even though the presence of an additional water molecule enhances forward CT, the overall net electron transfer remains negligibly small, and furthermore it accumulates on water not hexane.
@@ 2.2 Charge transfer at an extended oil-water interface
Analysis of minimized dimer and trimer configurations reveal that forward and backward CT between water and hexane nearly cancels. However, the CT mechanism at the extended water–oil interface may be significantly more intricate due to thermal fluctuations and a more extended network of hydrogen-bond interactions between waters and hydrophobic association of hexanes. To gain further insight into the CT process at an extended water–oil interface, molecular dynamics at ambient conditions using the polarizable AMOEBA force fieldponder2010current, ren2011polarizable was employed; simulation details are provided in the Methods section.
@@
From the MD simulations we draw two types of interfacial hexane-water clusters. Figure 2A is a representative snapshot of the 200 clusters in which the water molecules are localized around multiple hexanes, whereas Figure 2B is a representative snapshot of the 50 clusters in which interfacial water molecules form more extensive hydrogen-bonded network around the hexane. From these clusters we derive the probability density distribution in Figure 2C that shows a high-probability region near a CH…O angle of 120° for C-O distances >> 3.5 Å, consistent with the C-H···O configurations of the ab initio optimized clusters in Figure 1B,C. The broad distribution means that the C-H···O hydrogen bond is significantly weaker and more distorted compared to the highly directional hydrogen bonds found for thermalized water–water interactions which experience significant CT. This would indicate that CT in the forward direction from water to oil will be much smaller at an extended oil-water interface as well. In Figure 2D, the distribution of distances between hexane and water hydrogens is centered at 1.85 Å, whereas the distribution of water oxygen to hexane hydrogen distances is peaked at 2.2 Å, and the two distributions exhibit minimal overlap. This interfacial organization implies that the hydrogen of surface water orientated towards the oil (“dangling O-H groups”) will be a conduit for back-transfer of electron density from oil to water.
@@
This is borne out in Figure 2E which illustrates the ensemble nature of CT dynamics at the oil-water interface. The forward CT distribution of the interfacial clusters peaks at 2.09 me-/nm2, which arises from the C-H···O hydrogen bond at the interface, and similar to the small clusters in Figure 1. But again this is balanced by electron backflow from the CT distribution centered around -1.44 me-/nm2 due to the presence of short-range H···H interactions at the interface that defines the source of backward CT as also seen in Figure 1 The average net CT of 0.00065 e-/nm2 (0.65 me-/nm2) obtained from ALMO-EDA corresponds to 0.00007 e-/hydrogen bond, closely matching the 0.0001 e-/hydrogen bond
@@
Figure 2: Water/oil interface structures. (A, B) Sample representation of the clusters considered for the ALMO-EDA calculations highlighting the spatial arrangement of water molecules (red and white spheres) around hexane molecules (cyan). The extended hydrogen-bonded water network at the interface is illustrated in pink. (C) Probability density distribution illustrating the spatial and angular relationship between water oxygen and hexane. The x-axis represents the C–O distance, while the y-axis shows the C-H···O angle, describing the relative orientation of water molecules with respect to hexane’s C–H bonds. (D) Probability distribution of the closest distance between hydrogen atoms in hexane with water hydrogens and oxygens. This is consistent with the radial distribution functions (Supplementary Figure S1), where the H···H intermolecular distribution is more short-ranged than the O···H distribution in the water–hexane system. (E) Probability density distributions of Forward CT, Backward CT, and Net CT. Solid lines for direct water–hexane interactions clusters in (A) and dashed lines correspond to the clusters in (B). The original data is provided in Supplementary Tables S3 and S4. (F) Probability density distributions of water oil inter-molecule interaction energy components. The legend includes the percentage contribution of each energy component relative to the total interaction energy. Forward CT refers to H2O $\to$ C6H14, and Backward CT refers to C6H14 $\to$ H2O. Cluster selection details are provided in the Methods section.
@@
observed in the hexane–water clusters in Figure 1. Furthermore these conclusions are consistent between the clusters with localized waters near hexane or the clusters incorporating the extended hydrogen bonding network since interfacial CT is predominantly dictated by direct water–hexane interactions, whereas the broader hydrogen bonding network within the interfacial region exerts only a subtle perturbative influence (average net CT of 0.00061 e-/nm2). Finally Figure 2F shows that the frozen energy term constitutes the dominant component (61.6%) to water-hexane interactions, with similar percentages of polarization energy and CT making up the remaining energetic interactions. The estimated 20.3% of CT energy between water and hexane is a proportion similar to that observed in Figure 1 for small water–hexane clusters B and C, and is energetically the manifestation of electron delocalization in stabilizing the system.
2.3 Surface charge and oil-water emulsion metastability